# Student Member, IEEE

### 28 Feb 2024 v1

@@ Abstract
1 INTRODUCTION
@@
Thus, our approach facilitates efficient use of training data, informed classification decisions based on Bayesian principles, and enhanced robustness against increased uncertainty. Our method demonstrates a notable improvement in classification accuracy, with relative percentage increases of 35.71% and 14.14% at an SNR of 0 dB compared to single radar and RBC with soft-voting methods, respectively. Moreover, our method achieves a desired classification accuracy (or correct individual prediction) in a shorter dwell time compared to single radar and RBC with soft-voting methods
@@ 2 RELATED WORK
DL: The enhanced capability of radar to extract complex target features has sparked an increasing demand for models capable of using complex data, exemplified by the rising use of DL model [28, 1]. The work in [45] was the first application of Convolutional Neural Network (CNN) for monostatic radar classification based on simulated RCS time series data of geometric shapes. However, CNNs do not incorporate long-term temporal dynamics into their modeling approach. To address this limitation and capture long-term and short term temporal dynamics, many studies have used Recurrent Neural Networks (RNNs) and Long Short Term Memorys (LSTMs) networks [46]. These works classify a sliding time window or the entire trajectory of RCS time series data corresponding to monostatic radars tracking an UAV [47, 48, 49]. [50] combined a bidirectional Gated Recurrent Unit (GRU) RNN and a CNN to extract features from RCS time series data of geometric shapes and small-sized planes, subsequently using an Feed Forward Neural Network (FFNN) for classification. Lastly, [29] employed a spatio-temporal-frequency Graph Neural Network (GNN) with RNN aggregation to classify two simulated aircraft with different micromotions tracked by a synthetic heterogeneous distributed radar system. While numerous DL approaches have been utilized in RCS based RATR classification frameworks, challenges still persist, including the necessity for substantial data to generalize to unseen data [51], overconfident predictions [52], and a lack of interpretability [53]. Furthermore, Ezuma et al showed that Tree-based ML models outperformed traditional deep complex Neural Network (NN) architectures for commercial drone classification. Thus this paper focuses on the following ML models: Extreme Gradient Boosting (XGBoost), MLP with three hidden layers, and logistic regression.
3 Multi-Static RCS Dataset Generation
@@ 3.1 Existing Data
3.2 Training Data Generation
@@
where $x_{i}^{(t)}=[\sigma_{i}^{(t)},\phi_{i}^{(t)},\theta_{i}^{(t)}]$ contains the noisy RCS signature, the noisy azimuth and the noisy elevation. The noisy RCS signature, azimuth, and elevation will be discussed in Section 3.
@@
3.3 Testing Data Generation
@@
$\gamma^{(t)}$ $=\gamma^{(t-1)}+\epsilon_{\gamma};$ $\epsilon_{\beta}\sim\mathcal{N}(0^{\circ},144^{\circ})$ (4)
@@
We train the discriminative ML model on the RCS, azimuth, and elevation measurements described in Section 3 Subsection 3.2, and evaluate our RATR framework on the trajectories described in Section 3 Subsection 3.3. We next discuss how these discriminative ML models are integrated into the RATR framework.
4 METHODOLOGY
At each time step $t$ multiple radars pulse EMWs at the target UAV, where each radar $j$ calculates a RCS signature, azimuth, and elevation $x_{j}$. Each radar inputs its respective observation to a local discriminative ML model to output a UAV type probability vector. Here, local denontes that the ML resides at the individual radar hardware. All the individual radar UAV type probability vectors are combined by a fusion method (Section 4 Subsection 4.2). Subsequently, the posterior UAV type probability distribution, based on the past RCS signatures $z^{(1:t-1)}$, is updated with the fused UAV type probability distribution from time step $t$. A RBC framework is applied such that the posterior probability distribution recursively updates as the radars continuously track a moving UAV. Our framework is shown in Figure 4.
@@ 4.2 Fusion Methods
4.2.1 Bayesian Optimal Fusion
@@ 4.2.3 Hard-Voting
$P(C=c|Z^{(t)})=\begin{cases}1-\frac{|\mathcal{C}|\epsilon}{|\mathcal{C}|+1}&\text{if }c=Mo(\hat{c_{1}},\ldots,\hat{c_{J}})\\ \frac{\epsilon}{|\mathcal{C}|}&\text{otherwise}\end{cases}$ (19)
@@ 4.2.5 Maximum
4.3 Recursive Bayesian Classification
@@ 5.1 Configuration
5.2 Signal-to-Noise Ratio
@@ 5.3 Results
We visualize the RBC accuracy as the SNR changes (Figure 7). We observe an ”S” curve by following the top of the barplot across columnns of figure axes, where an increase in the signal power does not lead to an increase in classification performance at the upper extremes of SNR. However, radar channels typically operate in the 0 to -20 dB SNR range. When the RATR is operating at lower SNR ranges, having more observations of the target UAV substantially improves the classification performance versus a single radar (Figure 6). If the SNR is high, the classification performance of a single radar should approach that of multi-radar classification, given a sufficiently large dwell time and suitable target geometries, especially when using a complex ML model (Figure 8).

### 08 Mar 2024 v2

# Multistatic-Radar RCS-Signature Recognition of Aerial Vehicles: A Bayesian Fusion Approach

@@ Abstract
Acknowledgements
Research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-23-2-0014. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
I INTRODUCTION
@@
Thus, our approach facilitates efficient use of training data, informed classification decisions based on Bayesian principles, and enhanced robustness against increased uncertainty. Our method demonstrates a notable improvement in classification accuracy, with relative percentage increases of 35.71% and 14.14% at an SNR of 0 dB compared to single radar and RBC with soft-voting methods, respectively. Moreover, our method achieves a desired classification accuracy (or correct individual prediction) in a shorter dwell time compared to single radar and RBC with soft-voting methods.
@@ II RELATED WORK
DL: The enhanced capability of radar to extract complex target features has sparked an increasing demand for models capable of using complex data, exemplified by the rising use of DL model [28, 1]. The work in [44] was the first application of Convolutional Neural Network (CNN) for monostatic radar classification based on simulated RCS time series data of geometric shapes. However, CNNs do not incorporate long-term temporal dynamics into their modeling approach. To address this limitation and capture long-term and short term temporal dynamics, many studies have used Recurrent Neural Networks (RNNs) and Long Short Term Memorys (LSTMs) networks [45]. These works classify a sliding time window or the entire trajectory of RCS time series data corresponding to monostatic radars tracking an UAV [46, 47, 48]. [49] combined a bidirectional Gated Recurrent Unit (GRU) RNN and a CNN to extract features from RCS time series data of geometric shapes and small-sized planes, subsequently using an Feed Forward Neural Network (FFNN) for classification. Lastly, [29] employed a spatio-temporal-frequency Graph Neural Network (GNN) with RNN aggregation to classify two simulated aircraft with different micromotions tracked by a synthetic heterogeneous distributed radar system. While numerous DL approaches have been utilized in RCS based RATR classification frameworks, challenges still persist, including the necessity for substantial data to generalize to unseen data [50], overconfident predictions [51], and a lack of interpretability [52]. Furthermore, Ezuma et al showed that Tree-based ML models outperformed traditional deep complex Neural Network (NN) architectures for commercial drone classification. Furthermore [53] showed that for tabular datasets that DL does not perform as well as ML. Thus this paper focuses on the following ML models: Extreme Gradient Boosting (XGBoost), MLP with three hidden layers, and logistic regression.
III Multi-Static RCS Dataset Generation
@@ III-A Existing Data
III-B Training Data Generation
@@
where $x_{i}^{(t)}=[\sigma_{i}^{(t)},\phi_{i}^{(t)},\theta_{i}^{(t)}]$ contains the noisy RCS signature, the noisy azimuth and the noisy elevation. The noisy RCS signature, azimuth, and elevation will be discussed in Section III.
@@
III-C Testing Data Generation
@@
$\gamma^{(t)}$ $=\gamma^{(t-1)}+\epsilon_{\gamma};$ $\epsilon_{\gamma}\sim\mathcal{N}(0^{\circ},144^{\circ})$ (4)
@@
We train the discriminative ML model on the RCS, azimuth, and elevation measurements described in Section III Subsection III-B, and evaluate our RATR framework on the trajectories described in Section III Subsection III-C. We next discuss how these discriminative ML models are integrated into the RATR framework.
IV METHODOLOGY
At each time step $t$ multiple radars pulse EMWs at the target UAV, where each radar $j$ calculates a RCS signature, azimuth, and elevation $x_{j}$. Each radar inputs its respective observation to a local discriminative ML model to output a UAV type probability vector. Here, local denontes that the ML resides at the individual radar hardware. All the individual radar UAV type probability vectors are combined by a fusion method (Section IV Subsection IV-B). Subsequently, the posterior UAV type probability distribution, based on the past RCS signatures $z^{(1:t-1)}$, is updated with the fused UAV type probability distribution from time step $t$. A RBC framework is applied such that the posterior probability distribution recursively updates as the radars continuously track a moving UAV. Our framework is shown in Figure 4.
@@ IV-B Fusion Methods
IV-B1 Bayesian Optimal Fusion
@@ IV-B3 Hard-Voting
$P(C=c|Z^{(t)})=\begin{cases}1-\frac{(|\mathcal{C}|-1)\epsilon}{|\mathcal{C}|}&\text{if }c=Mo(\hat{c_{1}},\ldots,\hat{c_{J}})\\ \frac{\epsilon}{|\mathcal{C}|}&\text{otherwise}\end{cases}$ (19)
@@ IV-B5 Maximum
IV-C Recursive Bayesian Classification
@@ V-A Configuration
V-B Signal-to-Noise Ratio
@@ V-C Results
We visualize the RBC accuracy as the SNR changes (Figure 7). We observe an ”S” curve by following the top of the barplot across columnns of figure axes, where an increase in the signal power does not lead to an increase in classification performance at the upper extremes of SNR. However, radar channels typically operate in the 0 to -20 dB SNR range. When the RATR is operating at lower SNR ranges, having more observations of the target UAV substantially improves the classification performance versus a single radar (Figure 8).
If the SNR is high, the classification performance of a single radar should approach that of multi-radar classification, given a sufficiently large dwell time and suitable target geometries, especially when using a complex ML model (Figure 6).